OpenStudy (anonymous):

A line passing through (5, 6) and ( -1, -1/2)

OpenStudy (addonexus):

Anything else?

jigglypuff314 (jigglypuff314):

\[slope = m = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given the points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

i have to write my answer in standard form which is Ax+By =C

jigglypuff314 (jigglypuff314):

okay, so you would want to convert what you were given to a form/equation you know were you taught point-slope form or slope-intercept form?

OpenStudy (anonymous):

isn't point slope y-y1=m(x-x1) @jigglypuff314

jigglypuff314 (jigglypuff314):

yep :) okay so first you would want to find the slope between the given points

OpenStudy (anonymous):

okay when i subtract -1/2 -6 it equals -6 1/2 @jigglypuff314 that can't be right

jigglypuff314 (jigglypuff314):

that is what I got tho.

OpenStudy (anonymous):

how can that be a point ? @jigglypuff314

OpenStudy (anonymous):

how can you divide -6 1/2 over -6

jigglypuff314 (jigglypuff314):

it's not a point... it's the rate of change. -1/2 - 6 = -13/2 - 1 - 5 = -6 to simplify... \[\frac{ \frac{ -13 }{ 2 } }{ -6 } \times \frac{ 2 }{ 2 } = \frac{ -13 }{ -12 } = \frac{ 13 }{ 12 }\] so 13/12 is your slope

OpenStudy (anonymous):

how would you make that into Standard form @jigglypuff314

jigglypuff314 (jigglypuff314):

so first what did you get as your equation in point slope form?

OpenStudy (anonymous):

wait its suppose to be in point slope... i did it slope rise/run

OpenStudy (anonymous):

nevermind i get it

OpenStudy (anonymous):

how would you simpfiy 5 5/12? @jigglypuff314

jigglypuff314 (jigglypuff314):

\[5 \frac{ 5 }{ 12 } = \frac{ 12*5 }{ 12 } + \frac{ 5 }{ 12 } = \frac{ 60 }{ 12 } + \frac{ 5 }{ 12 } = \frac{ 65 }{ 12 }\]

OpenStudy (anonymous):

\[y-6=\frac{ 13 }{ 12 }x-\frac{ 65 }{ 12 }\] what now? @jigglypuff314

jigglypuff314 (jigglypuff314):

multiply all values by 12 because standard form doesn't want fractions/deacimals

OpenStudy (anonymous):

do i multiple everything by 12 to get rid of the fraction?

jigglypuff314 (jigglypuff314):

lol yep :)

OpenStudy (anonymous):

would the answer be 12y=13x-132 @jigglypuff314

jigglypuff314 (jigglypuff314):

once you multiply all by 12 you would get 12y - 72 = 13x - 65 then you would add 72 to both sides +72 +72 12y = 13x + 7 then standard form is ax + by = c so get x to the other side 13x - 12y = -7 is your answer :)

OpenStudy (anonymous):

OH darn i was close tho :-) @jigglypuff314 thank you tho

jigglypuff314 (jigglypuff314):

your welcome :D